OpenStudy (ammarah):

Classify the conic and write the equation in standard form. 8x^2-25y+32x+150y-393=0 I know that its a hyperbola so we plug it in for that equation...

OpenStudy (ammarah):

@ganeshie8

ganeshie8 (ganeshie8):

that would be a mistake

ganeshie8 (ganeshie8):

its not an hyperbola

ganeshie8 (ganeshie8):

hyperbola will have both \(\large x^2\) and \(\large y^2\) terms, right ?

ganeshie8 (ganeshie8):

look at the equation, u dont have any \(\large y^2\) terms

OpenStudy (ammarah):

im sorry its y^2 also

OpenStudy (ammarah):

8x^2-25y^2+32x+150y-393=0

ganeshie8 (ganeshie8):

Ohk... then you're right :) its a hyperbola !

ganeshie8 (ganeshie8):

u knw the standard form right ?

OpenStudy (ammarah):

yes

ganeshie8 (ganeshie8):

\(\large \dfrac{(x-h)^2}{a^2} - \dfrac{(y-k)^2}{b^2} = 1\)

ganeshie8 (ganeshie8):

its going to be very easy... u just need to complete the squre....

ganeshie8 (ganeshie8):

step1 : group x terms, group y terms, and slap the constant to the other side

ganeshie8 (ganeshie8):

\(\large 8x^2-25y^2+32x+150y-393=0\) group same variable terms : \(\large 8x^2+32x-25y^2 + 150y = 393\)

OpenStudy (ammarah):

ok i got 8(x^2+4) - 25(y^2+150y)

ganeshie8 (ganeshie8):

looks good, except for mistake in sign : you should get : \(\large 8x^2-25y^2+32x+150y-393=0\) group same variable terms : \(\large 8x^2+32x-25y^2 + 150y = 393\) \(\large 8(x^2+4x)-25(y^2 - 6y) = 393\)

OpenStudy (ammarah):

yes

ganeshie8 (ganeshie8):

next complete the square for the stuff inside parenthesis by adding and subtracting the square term : \(\large 8x^2-25y^2+32x+150y-393=0\) group same variable terms : \(\large 8x^2+32x-25y^2 + 150y = 393\) \(\large 8(x^2+4x)-25(y^2 - 6y) = 393\) \(\large 8(x^2+4x + 2^2 - 2^2)-25(y^2 - 6y + 3^2 -3^2) = 393\) complete the square : \(\large 8((x+2)^2 - 2^2)-25((y-3)^2 -3^2) = 393\)

ganeshie8 (ganeshie8):

see if that still looks fine :)

OpenStudy (ammarah):

yes an now we add the squares to the other side?

OpenStudy (ammarah):

oh wait noo its a hyperbola

ganeshie8 (ganeshie8):

exactly ! \(\large 8x^2-25y^2+32x+150y-393=0\) group same variable terms : \(\large 8x^2+32x-25y^2 + 150y = 393\) \(\large 8(x^2+4x)-25(y^2 - 6y) = 393\) \(\large 8(x^2+4x + 2^2 - 2^2)-25(y^2 - 6y + 3^2 -3^2) = 393\) complete the square : \(\large 8((x+2)^2 - 2^2)-25((y-3)^2 -3^2) = 393\) \(\large 8(x+2)^2 - 8*2^2-25(y-3)^2 +25*3^2 = 393\) \(\large 8(x+2)^2 - 32-25(y-3)^2 + 225 = 393\) \(\large 8(x+2)^2 -25(y-3)^2 + 193= 393\) \(\large 8(x+2)^2 -25(y-3)^2 = 200\)

ganeshie8 (ganeshie8):

see if you're fine so far

ganeshie8 (ganeshie8):

lastly, divide 200 both sides

ganeshie8 (ganeshie8):

\(\large 8(x+2)^2 -25(y-3)^2 = 200\) divide 200 both sides : \(\large \dfrac{ 8(x+2)^2}{200} -\dfrac{25(y-3)^2}{200} = 1\) simplify : \(\large \dfrac{ (x+2)^2}{25} -\dfrac{(y-3)^2}{8} = 1\)

ganeshie8 (ganeshie8):

we're done.

OpenStudy (ammarah):

ok yess im sorry to trouble u more but im stuck with writing equations of the graphs...

ganeshie8 (ganeshie8):

np :) i have time... go ahead